# An exact analytical solution for the weakly magnetized flow around an axially symmetric paraboloid, with application to magnetosphere models

### v4

@@ 4.1 Different aspect ratios
$\vec{B}_{\mathrm{tr}}^{\mathrm{sc}}(\vec{r}\,):=$ $\ B_{\mathrm{tr},\rho}(\vec{r}_{\ell})\,\vec{e}_{\rho}+B_{\mathrm{tr},\varphi}(\vec{r}_{\ell})\,\vec{e}_{\varphi}+$
@@
It seems worth noting that, while Equations (46)–(48), as well as Equation (52) together with its associated magnetic field (see Equations (B11)–(B13) in Appendix B) do constitute valid MHD solutions, the scaled velocity fields (46) and (52) are irrotational only for $\ell_{\rho}=\ell_{z}$ and $\ell_{x}=\ell_{y}=\ell_{z}$, respectively, and therefore in general do not have scalar potentials. The exact potential flows around scaled paraboloids could probably only be found be solving the Laplace equation (3) for these geometries, which is beyond the scope of this paper.
@@ 4.2 Advecting arbitrary magnetic fields
For the computation of the poloidal components, however, we expand the first and third components of Equation (55) up to first order
@@
independently of the specific form of $a$ or $\Psi$ (and hence independently of the shape of the obstacle). Then inserting these expressions into Equation (4.2) and solving for the desired poloidal field components (54) results in
@@
Here, the shape of the obstacle enters through the partial derivatives of ${\beta_{\mathrm{ren}}}$, which are already known from Equation (3.2) because $\beta$ and ${\beta_{\mathrm{ren}}}$ share the same gradient by construction. Inserting these into Equation (66) together with the components (14) and (15) of $\vec{u}$ and performing a series of algebraic simplifications, the explicit forms of the two poloidal components finally becomes
@@ 4.3 A simple example: cylindrical flux ropes
which has been used as a simple model for solar or interplanetary flux tubes (e.g. Gold & Hoyle, 1960; Kleimann & Hornig, 2001; Aschwanden, 2019). Its field lines form helical spirals of constant axial radius and constant thread lead $2\pi$. To make this example field more interesting, we reduce its symmetry by first rotating it by $90^{\circ}$ about the $y$-axis such that the $x$-axis becomes the new axis of symmetry, and then displace this axis of symmetry from $(y,z)=(0,0)$ to $(y_{\mathrm{c}},z_{\mathrm{c}})$. The result reads
$\vec{B}_{0}=$ $\ \bigl{(}\left[\cos\varphi-(z-z_{\mathrm{c}})\sin\varphi\right]\vec{e}_{\rho}\bigr{.}-$
$\ \left[\sin\varphi+(z-z_{\mathrm{c}})\cos\varphi\right]\vec{e}_{\varphi}+\big{.}\left[\rho\sin\varphi-y_{\mathrm{c}}\right]\vec{e}_{z}\bigr{)}\times$
$\ \bigl{[}1+(\rho\sin\varphi-y_{\mathrm{c}})^{2}+(z-z_{\mathrm{c}})^{2}\bigr{]}^{-1}.$ (71)
@@
Figure 3: Left: perspective rendering of several instances of the sample flux tube field (4.3) with its axis parallel to the $x$-axis and passing through $y_{\mathrm{c}}=\pm 2$ and $z_{\mathrm{c}}\in\{-3,0,3\}$, represented in its original form ($y<0$) vs. its shape when distorted and displaced by the presence of the parabolic obstacle ($y>0$). A semi-transparent cylinder of radius $0.8$ marks the outer boundary of the flux tube in each case. The color of field lines encodes the total field strength, increasing from blue to red. Middle: orthogonal projection of the same situation, viewed from the $+x$ direction. Right: again the same situation viewed from $+y$ direction. Note that, since this is also an orthogonal projection, the deformed flux tube segments in the front are indeed physically longer than those on the far side, rather than just appearing so because of their closer proximity to the observer.
@@ 4.4 Advecting arbitrary scalar fields
Figure 4: The deformation of basic geometric shapes induced by the presence of the parabolic obstacle. (A): sequence of red spheres of radius $0.8$ centered at $(x_{\mathrm{c}},y_{\mathrm{c}},z_{\mathrm{c}})=(-1,-1,t_{z})$ for $t_{z}\in\{-3,-1,1,3\}$, together with a sequence of blue cubes of side length $1.2$ centered at $(x_{\mathrm{c}},y_{\mathrm{c}},z_{\mathrm{c}})=(1,1,t_{z})$ and arranged parallel to, but not intersecting, the inflow axis. (B): The same shapes are severely distorted when passively advected in the flow, which is visualized by the pancake-like compression and flattening of these shapes as they are advected towards the obstacle, eventually sliding off it. (C): The same situation as in (A) but for a sequence of cylinders of height $1.6$ and radius $0.8$ alternating in color and orientation and placed right on the $z$-axis at $(x_{\mathrm{c}},y_{\mathrm{c}},z_{\mathrm{c}})=(0,0,t_{z})$. (D): The cylinders are being compressed into (eventually) infinitesimally thin shells coating the paraboloid’s surface.
@@ 4.5 The inverse transformation
as an implicit equation for the coordinate $z$. For fixed $a$, this equation only depends on a single argument, and may be interpreted as giving the isochrone label of a fluid element located at height $z$ while sliding along streamline $a$. It is clear from this interpretation, and also easy to verify via
@@
that $z\mapsto\beta_{a}(z)$ is strictly monotonous. And since it is also strictly concave, i.e. $\beta_{a}^{\prime\prime}(z)<0$ for $z>1/2$, Equation (76) is well suited for an iterative solver based on Newton’s method, which in this case reads
@@
For the special case $a=0$, we need to bear in mind that, since this is the only streamline that does not extend indefinitely into the negative $z$ direction but rather ends at the stagnation point, it may happen that $z_{k+1}\leq 1/2$ for some $k$, indicating a position on or inside the paraboloid. A possible strategy, which we shall adopt here without examining further options for its optimization, is to replace $z_{k+1}$ by $(1/2+z_{k})/2$, i.e. the midpoint between the stagnation point and the previous approximate value. This ensures that $z_{k+1}$ is both a valid exterior position and closer to the desired result than $z_{k}$ had been.
@@ 5.1 Density and flow fields
to be satisfied. Here $\gamma$ is the adiabatic index and the new parameter $m\in[0,1]$ denotes the sonic Mach number at upstream infinity, with $P$ consistently being normalized to its respective value at the upstream boundary.
@@
Although the roots of the quartic (86) can in principle be found using Ferrari’s formula, a more convenient expression would be desirable, and the same is certainly true for Equation (85) involving the Lambert $W$ function. For this reason, we again repeat the same steps already applied in Kleimann et al. (2017), namely to replace both Equation (85) and $n({\cal A})$ in the exact (but implicit) expression (86) by the second-order polynomial $n_{2}({\cal A})$ whose coefficients are adjusted such that it coincides with $n({\cal A})$ at ${\cal A}=0$ (the stagnation point) and ${\cal A}=1$ (infinity), and also has the same derivative at ${\cal A}=0$. As a simpler, less accurate alternative, we also consider the first-order polynomial $n_{1}({\cal A})$, which only fulfills the first two conditions. Reproducing Equations (44)–(47) from Kleimann et al. (2017), we get
@@
It is worth noting that, while the polynomial approximation can be evaluated at any position $(\rho,z)$, the exact expression exhibits a region in which it becomes complex-valued, and hence meaningless (indicated in white in Figure 5). In the heliosphere case, this region grows as $m$ increases, eventually engulfing the entire downwind half-space as $m\rightarrow 1$. This, however, is apparently different for the paraboloid, where it remains confined to the interior of the obstacle, thereby allowing the exact expression (85) to be used for any $m\leq 1$ (and technically even beyond that, though of course no bow shock can be expected) if so desired.
@@ 5.2 The magnetic field for compressible flow
$\ (\omega/n)\,\underbrace{\nabla\cdot\vec{u}^{\mathrm{I}}}_{=0}+\left(\vec{u}^{\mathrm{I}}\cdot\nabla\right)(\omega/n)$ (91)
@@
where $x=\rho\,\cos\varphi$, $\varphi\in\{0,\pi\}$, is used rather than $\rho$ to be able to access the negative values of the $x$ coordinate. This method allows us to draw field lines without the otherwise mandatory recourse to a numerical field line tracer, as exemplified in Figure 6. Therein, contours of expression (104) are drawn using the compressible second Euler potential (D), evaluated for the first-order density approximation (87), i.e. with $\nu_{1}=1-n_{\mathrm{sp}}$ and $\nu_{2}=0$. With the corresponding field lines (derived by traditional numerical integration) overplotted, and both adjusted to pass through the same set of seed points, the excellent agreement of the results from both methods is immediately evident.
@@ 6 Summary and conclusions
In this work, we presented a (magneto)hydrodynamical model for laminar, inviscid flow ahead of and around a solid obstacle of paraboloid form whose axis of symmetry coincides with that of the incident flow. For the case of the fluid being both incompressible and irrotational, we derived compact formulas for the scalar flow potential and the stream function, and hence for the velocity field. This velocity field was then used as an input to the induction equation of ideal magnetohydrodynamics, which we solved for the magnetic field subject to a homogeneous upstream boundary condition of arbitrary inclination. The resulting field components can be written in compact and explicit form, involving neither integrals nor special functions. This model constitutes the first (and therefore so far also the only) known nontrivial axially symmetric magnetohydrodynamical solution for obstacle-modulated flow with this property.
@@
In a further step, we generalized the above model first from circular to elliptic paraboloids and then from incompressible to mildly compressible (subsonic) flow. The resulting spatial variations in density, which exhibit the expected pile-up ahead of the stagnation point and a return to the undisturbed density at large lateral distances, are consistent with conservation of mass and approximately consistent with conservation of linear momentum along streamlines. Even for this extension to compressibility, we found an exact solution of the associated magnetic field frozen into the flow at the expense of more involved, albeit still fully analytical expressions without recourse to special functions or unevaluated integrals. Already in the simpler incompressible case, this property is unique among known analytical solutions for magnetized flow around solid obstacles, and therefore makes the present model a valuable tool not only in the magnetospheric context, but also for qualitative investigations of similar geometries such as the sphere/ellipsoid or the heliopause.
@@ B.2 Cartesian coordinates
$B_{z}$ $=-\frac{\left[\left(x/\ell_{x}^{2}\right)B_{x0}+\left(y/\ell_{y}^{2}\right)B_{y0}\right]\ell_{z}^{2}}{2\,r_{\ell}\,\sqrt{\left(r_{\ell}\,\ell_{z}+z\right)\left([r_{\ell}-1]\ell_{z}+z\right)}}+\left(1-\frac{1}{2\,r_{\ell}}\right)B_{z0}\,,$ (B13)
@@ Appendix C Determination of the functions $\vec{K}_{1}$ and $\vec{K}_{2}$
We present explicit analytical expressions for the functions $\vec{K}_{1}$ and $\vec{K}_{2}$ defined in Equation (94). Although we provide the formula for the function $\vec{K}_{2}$ for the sake of completeness, for many applications it will likely be sufficient to work with the first-order density approximation (87), whose magnetic field merely requires the simpler $\vec{K}_{1}$ but not the much more involved $\vec{K}_{2}$ (see below). Besides its more complex formulas, a further potential complication with the second-order case may arise in numerical applications because several of the terms in $\vec{K}_{2}$ individually diverge in the vicinity of the $(a=0)$ streamline, and only produce values of order unity in their total sum through mutual cancellation. This is the main reason why the second-order magnetic field formulas have not been used in Figure 6.
@@
is a direct result of the condition $a(\rho_{0},z_{0})=a(\rho,z)$. Evaluating the gradients on the right hand side of (C1), we obtain
@@
$\frac{{\cal A}}{u_{\rho}^{\mathrm{I}}}=\rho+\frac{a^{4}}{(\rho^{2}-a^{2})^{2}\,\rho}$ (C4)
@@
the remaining integral on the right hand side for the case $k=1$ readily yields
@@
such that the integral on the right hand side of Equation (C3) for $k=2$ becomes
$\int_{\rho_{0}}^{\rho}\partial_{a}\biggl{(}\frac{{\cal A}^{2}}{u_{\rho}^{\mathrm{I}}}\biggr{)}\,{\rm d}\rho^{\prime}=$ $\ -\frac{2}{a}\,\biggl{(}\frac{1}{a^{4}}-\frac{4}{a^{2}}+2\biggr{)}\,Q-\frac{2\,(\rho^{2}-\rho_{0}^{2})\,\bigl{[}\rho^{2}\,\rho_{0}^{2}+(1-3\,a^{2})\,(\rho^{2}+\rho_{0}^{2})-5\,(1-a^{2})\,a^{2}+1\bigr{]}}{\bigl{[}\rho^{2}+(\rho^{2}-a^{2})^{2}\bigr{]}\,\bigl{[}\rho_{0}^{2}+(\rho_{0}^{2}-a^{2})^{2}\bigr{]}\,a}-$
@@
$\vec{K}_{2}=\frac{1}{2\,r}\,\sqrt{\frac{r+z}{r-z}}\,\biggl{[}2\,(r-1)+\frac{1}{r+z}\biggr{]}^{2}\,\vec{e}_{\rho}+\biggl{[}\biggl{(}1-\frac{1}{2\,r}\biggr{)}\,\sqrt{\frac{r+z}{r-z}}\ \vec{e}_{\rho}+\frac{1}{2\,r}\,\vec{e}_{z}\biggr{]}\times\biggl{[}2\,\biggl{(}1-\frac{\rho^{2}}{a^{2}}\biggr{)}\,\biggl{(}\frac{1}{a^{4}}-\frac{4}{a^{2}}+2\biggr{)}\times\biggr{.}$
.×$\biggl{.}\times\,Q(\rho,a;a)+3-2\,r-\frac{r^{2}+2\,z^{2}+3\,r\,z-4\,r-5\,z+4+1/(r-z)}{(r+z-1)^{2}\,r}+\frac{(2\,a^{2}-1)\,(r-z)}{a^{6}}\,\ln{\biggl{(}\frac{2\,(r+z-1)\,r}{(r+z)^{2}}\biggr{)}}\biggr{]}\,.\hskip 34.14322pt$ (C12)

### v5

@@ 4.1 Different aspect ratios
$\vec{B}_{\mathrm{tr}}^{\mathrm{sc}}(\vec{r}\,):=$ $\ B_{\mathrm{tr},\rho}(\vec{r}_{\ell})\,\vec{e}_{\rho}+B_{\mathrm{tr},\varphi}(\vec{r}_{\ell})\,\vec{e}_{\varphi}$
@@
It seems worth noting that, while Equations (46)–(48) and also Equation (52) together with its associated magnetic field (see Equations (B11)–(B13) in Appendix B) do constitute valid MHD solutions, the scaled velocity fields (46) and (52) are irrotational only for $\ell_{\rho}=\ell_{z}$ and $\ell_{x}=\ell_{y}=\ell_{z}$, respectively, and therefore in general do not have scalar potentials. The exact potential flows around scaled paraboloids could probably only be found be solving the Laplace equation (3) for these geometries, which is beyond the scope of this paper.
@@ 4.2 Advecting arbitrary magnetic fields
For the computation of the poloidal components, however, we expand the first and third components of Equation (55) up to first order,
@@
independently of the specific form of $a$ or $\Psi$ (and hence independently of the shape of the obstacle). Then, inserting these expressions into Equation (4.2) and solving for the desired poloidal field components (54) results in
@@
Here, the shape of the obstacle enters through the partial derivatives of ${\beta_{\mathrm{ren}}}$, which are already known from Equation (3.2) because $\beta$ and ${\beta_{\mathrm{ren}}}$ share the same gradient by construction. Inserting these into Equation (66) together with the components (14) and (15) of $\vec{u}$ and performing a series of algebraic simplifications, the explicit forms of the two poloidal components finally become
@@ 4.3 A simple example: Cylindrical flux ropes
which has been used as a simple model for solar or interplanetary flux tubes (e.g., Gold & Hoyle, 1960; Kleimann & Hornig, 2001; Aschwanden, 2019). Its field lines form helical spirals of constant axial radius and constant thread lead $2\pi$. To make this example field more interesting, we reduce its symmetry by first rotating it by $90^{\circ}$ about the $y$-axis such that the $x$-axis becomes the new axis of symmetry, and then displace this axis of symmetry from $(y,z)=(0,0)$ to $(y_{\mathrm{c}},z_{\mathrm{c}})$. The result reads
$\vec{B}_{0}=$ $\ \bigl{(}\left[\cos\varphi-(z-z_{\mathrm{c}})\sin\varphi\right]\vec{e}_{\rho}\bigr{.}$
$\ -\left[\sin\varphi+(z-z_{\mathrm{c}})\cos\varphi\right]\vec{e}_{\varphi}+\big{.}\left[\rho\sin\varphi-y_{\mathrm{c}}\right]\vec{e}_{z}\bigr{)}$
×$\ \times\bigl{[}1+(\rho\sin\varphi-y_{\mathrm{c}})^{2}+(z-z_{\mathrm{c}})^{2}\bigr{]}^{-1}.$ (71)
@@
Figure 3: Left: perspective rendering of several instances of the sample flux tube field (4.3) with its axis parallel to the $x$-axis and passing through $y_{\mathrm{c}}=\pm 2$ and $z_{\mathrm{c}}\in\{-3,0,3\}$, represented in its original form ($y<0$) vs. its shape when distorted and displaced by the presence of the parabolic obstacle ($y>0$). A semi-transparent cylinder of radius $0.8$ marks the outer boundary of the flux tube in each case. The color of field lines encodes the total field strength, increasing from blue to red. Middle: orthogonal projection of the same situation, viewed from the $+x$-direction. Right: again the same situation viewed from $+y$-direction. Note that, since this is also an orthogonal projection, the deformed flux tube segments in the front are indeed physically longer than those on the far side, rather than just appearing so because of their closer proximity to the observer.
@@ 4.4 Advecting arbitrary scalar fields
Figure 4: The deformation of basic geometric shapes induced by the presence of the parabolic obstacle. (a) Sequence of red spheres of radius $0.8$ centered at $(x_{\mathrm{c}},y_{\mathrm{c}},z_{\mathrm{c}})=(-1,-1,t_{z})$ for $t_{z}\in\{-3,-1,1,3\}$, together with a sequence of blue cubes of side length $1.2$ centered at $(x_{\mathrm{c}},y_{\mathrm{c}},z_{\mathrm{c}})=(1,1,t_{z})$ and arranged parallel to, but not intersecting, the inflow axis. (b) The same shapes are severely distorted when passively advected in the flow, which is visualized by the pancake-like compression and flattening of these shapes as they are advected toward the obstacle, eventually sliding off it. (c) The same situation as in (a) but for a sequence of cylinders of height $1.6$ and radius $0.8$ alternating in color and orientation and placed right on the $z$-axis at $(x_{\mathrm{c}},y_{\mathrm{c}},z_{\mathrm{c}})=(0,0,t_{z})$. (d) The cylinders are being compressed into (eventually) infinitesimally thin shells coating the paraboloid’s surface.
@@ 4.5 The inverse transformation
as an implicit equation for the coordinate $z$. For fixed $a$, this equation only depends on a single argument and may be interpreted as giving the isochrone label of a fluid element located at height $z$ while sliding along streamline $a$. It is clear from this interpretation, and also easy to verify via
@@
that $z\mapsto\beta_{a}(z)$ is strictly monotonous. Since it is also strictly concave, i.e., $\beta_{a}^{\prime\prime}(z)<0$ for $z>1/2$, Equation (76) is well suited for an iterative solver based on Newton’s method, which in this case reads
@@
For the special case $a=0$, we need to bear in mind that, since this is the only streamline that does not extend indefinitely into the negative $z$ direction but rather ends at the stagnation point, it may happen that $z_{k+1}\leq 1/2$ for some $k$, indicating a position on or inside the paraboloid. A possible strategy, which we shall adopt here without examining further options for its optimization, is to replace $z_{k+1}$ by $(1/2+z_{k})/2$, i.e., the midpoint between the stagnation point and the previous approximate value. This ensures that $z_{k+1}$ is both a valid exterior position and closer to the desired result than $z_{k}$ had been.
@@ 5.1 Density and flow fields
to be satisfied. Here, $\gamma$ is the adiabatic index and the new parameter $m\in[0,1]$ denotes the sonic Mach number at upstream infinity, with $P$ consistently being normalized to its respective value at the upstream boundary. We note that the procedure to be described is also applicable to the more general equation of state $P=p_{0}+p_{1}\,n^{\gamma}$ with constants $p_{0}$ and $p_{1}>0$, which has been suggested in the context of suprathermal particle populations, particularly those described by $\kappa$-distributions (e.g., Silveira et al., 2021; Lazar et al., 2020, and references therein). In that case, the resulting density pattern does not depend on the “vacuum pressure” contribution $p_{0}$ because it is only the gradient of $P$ that enters Equation (80). The prefactor $p_{1}$, on the other hand, can conveniently be absorbed into a redefined Mach number $m^{\prime}=m/\sqrt{p_{1}}$. We further note that our derivation is not conditional on the interpretation of $\gamma$ as the fluid’s adiabatic index, but is just as valid for other constant exponents (e.g., Scudder, 1992; Meyer-Vernet et al., 1995). For these reasons, we will continue to use the Maxwellian adiabatic equation of state (81) in the ensuing considerations.
@@
Although the roots of the quartic (86) can in principle be found using Ferrari’s formula, a more convenient expression would be desirable, and the same is certainly true for Equation (85) involving the Lambert $W$ function. For this reason, we again repeat the same steps already applied in Kleimann et al. (2017), namely, to replace both Equation (85) and $n({\cal A})$ in the exact (but implicit) expression (86) by the second-order polynomial $n_{2}({\cal A})$ whose coefficients are adjusted such that it coincides with $n({\cal A})$ at ${\cal A}=0$ (the stagnation point) and ${\cal A}=1$ (infinity), and also has the same derivative at ${\cal A}=0$. As a simpler, less accurate alternative, we also consider the first-order polynomial $n_{1}({\cal A})$, which only fulfills the first two conditions. Reproducing Equations (44)–(47) from Kleimann et al. (2017), we get
@@
It is worth noting that, while the polynomial approximation can be evaluated at any position $(\rho,z)$, the exact expression exhibits a region in which it becomes complex-valued and hence meaningless (indicated in white in Figure 5). In the heliosphere case, this region grows as $m$ increases, eventually engulfing the entire downwind half-space as $m\rightarrow 1$. This, however, is apparently different for the paraboloid, where it remains confined to the interior of the obstacle, thereby allowing the exact expression (85) to be used for any $m\leq 1$ (and technically even beyond that, though of course no bow shock can be expected) if so desired.
@@ 5.2 The magnetic field for compressible flow
$=$ $\ (\omega/n)\,\underbrace{\nabla\cdot\vec{u}^{\mathrm{I}}}_{=0}+\left(\vec{u}^{\mathrm{I}}\cdot\nabla\right)(\omega/n)$ (91)
@@
where $x=\rho\,\cos\varphi$, $\varphi\in\{0,\pi\}$, is used rather than $\rho$ to be able to access the negative values of the $x$ coordinate. This method allows us to draw field lines without the otherwise mandatory recourse to a numerical field line tracer, as exemplified in Figure 6. Therein, contours of expression (104) are drawn using the compressible second Euler potential (D), evaluated for the first-order density approximation (87), i.e., with $\nu_{1}=1-n_{\mathrm{sp}}$ and $\nu_{2}=0$. With the corresponding field lines (derived by traditional numerical integration) overplotted, and both adjusted to pass through the same set of seed points, the excellent agreement of the results from both methods is immediately evident.
@@ 6 Summary and conclusions
In this work, we presented a (magneto)hydrodynamical model for laminar, inviscid flow ahead of and around a solid obstacle of paraboloid form whose axis of symmetry coincides with that of the incident flow. For the case of the fluid being both incompressible and irrotational, we derived compact formulas for the scalar flow potential and the stream function, and hence for the velocity field. This velocity field was then used as an input to the induction equation of ideal magnetohydrodynamics, which we solved for the magnetic field subject to a homogeneous upstream boundary condition of arbitrary inclination. The resulting field components can be written in compact and explicit form, involving neither unevaluated integrals nor special functions. This model constitutes the first (and therefore so far also the only) known nontrivial axially symmetric magnetohydrodynamical solution for obstacle-modulated flow with this property.
@@
In a further step, we generalized the above model first from circular to elliptic paraboloids and then from incompressible to mildly compressible (subsonic) flow. The resulting spatial variations in density, which exhibit the expected pileup ahead of the stagnation point and a return to the undisturbed density at large lateral distances, are consistent with conservation of mass and approximately consistent with conservation of linear momentum along streamlines. Even for this extension to compressibility, we found an exact solution of the associated magnetic field frozen into the flow at the expense of more involved, albeit still fully analytical expressions without recourse to special functions or unevaluated integrals. Already in the simpler incompressible case, this property is unique among known analytical solutions for magnetized flow around solid obstacles, and therefore makes the present model a valuable tool not only in the magnetospheric context, but also for qualitative investigations of similar geometries such as the sphere/ellipsoid or the heliopause.
@@ B.2 Cartesian coordinates
$B_{z}$ $=-\frac{\left[\left(x/\ell_{x}^{2}\right)B_{x0}+\left(y/\ell_{y}^{2}\right)B_{y0}\right]\ell_{z}^{2}}{2\,r_{\ell}\,\sqrt{\left(r_{\ell}\,\ell_{z}+z\right)\left([r_{\ell}-1]\ell_{z}+z\right)}}+\left(1-\frac{1}{2\,r_{\ell}}\right)B_{z0}$ (B13)
@@ Appendix C Determination of the functions $\vec{K}_{1}$ and $\vec{K}_{2}$
We present explicit analytical expressions for the functions $\vec{K}_{1}$ and $\vec{K}_{2}$ defined in Equation (94). Although we provide the formula for the function $\vec{K}_{2}$ for the sake of completeness, for many applications it will likely be sufficient to work with the first-order density approximation (87), whose magnetic field merely requires the simpler $\vec{K}_{1}$ but not the much more involved $\vec{K}_{2}$ (see below). In addition to its more complex formulas, a further potential complication with the second-order case may arise in numerical applications because several of the terms in $\vec{K}_{2}$ individually diverge in the vicinity of the $(a=0)$ streamline, and only produce values of order unity in their total sum through mutual cancellation. This is the main reason why the second-order magnetic field formulas have not been used in Figure 6.
@@
is a direct result of the condition $a(\rho_{0},z_{0})=a(\rho,z)$. Evaluating the gradients on the right-hand side of (C1), we obtain
@@
$\frac{{\cal A}}{u_{\rho}^{\mathrm{I}}}=\rho+\frac{a^{4}}{(\rho^{2}-a^{2})^{2}\,\rho}\,,$ (C4)
@@
the remaining integral on the right-hand side for the case $k=1$ readily yields
@@
such that the integral on the right-hand side of Equation (C3) for $k=2$ becomes
$\int_{\rho_{0}}^{\rho}\partial_{a}\biggl{(}\frac{{\cal A}^{2}}{u_{\rho}^{\mathrm{I}}}\biggr{)}\,{\rm d}\rho^{\prime}=$ $\ -\frac{2}{a}\,\biggl{(}\frac{1}{a^{4}}-\frac{4}{a^{2}}+2\biggr{)}\,Q-\frac{2\,(\rho^{2}-\rho_{0}^{2})\,\bigl{[}\rho^{2}\,\rho_{0}^{2}+(1-3\,a^{2})\,(\rho^{2}+\rho_{0}^{2})-5\,(1-a^{2})\,a^{2}+1\bigr{]}}{\bigl{[}\rho^{2}+(\rho^{2}-a^{2})^{2}\bigr{]}\,\bigl{[}\rho_{0}^{2}+(\rho_{0}^{2}-a^{2})^{2}\bigr{]}\,a}$
@@
$\vec{K}_{2}=$ $\frac{1}{2\,r}\,\sqrt{\frac{r+z}{r-z}}\,\biggl{[}2\,(r-1)+\frac{1}{r+z}\biggr{]}^{2}\,\vec{e}_{\rho}+\biggl{[}\biggl{(}1-\frac{1}{2\,r}\biggr{)}\,\sqrt{\frac{r+z}{r-z}}\ \vec{e}_{\rho}+\frac{1}{2\,r}\,\vec{e}_{z}\biggr{]}\times\biggl{[}2\,\biggl{(}1-\frac{\rho^{2}}{a^{2}}\biggr{)}\,\biggl{(}\frac{1}{a^{4}}-\frac{4}{a^{2}}+2\biggr{)}\biggr{.}Q(\rho,a;a)$
.$\biggl{.}+3-2\,r-\frac{r^{2}+2\,z^{2}+3\,r\,z-4\,r-5\,z+4+1/(r-z)}{(r+z-1)^{2}\,r}+\frac{(2\,a^{2}-1)\,(r-z)}{a^{6}}\,\ln{\biggl{(}\frac{2\,(r+z-1)\,r}{(r+z)^{2}}\biggr{)}}\biggr{]}\,.$ (C12)